OpenStudy (anonymous):

On a test, when a student gets full credit for an essay and answers 13 questions correctly, the student’s test score is 77. When a student gets full credit for an essay and answers18 questions correctly, the student’s test score is 97. Assume the relationship is linear. A line models the relationship between the number of questions correctly answered, x, and the test score, y. What is the slope of the line - I cannot seem to get this one.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

To find the slope between two points, use the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\). Here your two points are (13, 77) and (18, 97).

OpenStudy (igreen):

(13, 77) and (18, 97) \(y_2 = 97\) \(y_1 = 77\) \(x_2 = 18\) \(x_1 = 13\) Plug these terms in: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{97-77}{18-13}\) Subtract: \(m = \dfrac{20}{5}\) \(m = 4\) So the slope is 4.